OpenStudy (anonymous):

A box has 8 beads of the same size, but all are different colors. Tamy draws a bead randomly from the box, notes its color, and then puts the bead back in the box. She repeats this 3 times. What is the probability that Tamy would pick a red bead on the first draw, then a white bead, and finally a red bead again? 1 over 4096 1 over 512 1 over 64 1 over 8

OpenStudy (misty1212):

she puts them back each time right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

and the beads are all different colors, so 8 different colors the probability of any one color is therefore \(\frac{1}{8}\)

OpenStudy (misty1212):

that is not the final answer though

OpenStudy (anonymous):

oh ok

OpenStudy (misty1212):

do it 3 times in a row, the probability of red, white, red is\[\frac{1}{8}\times \frac{1}{8}\times \frac{1}{8}\]

OpenStudy (misty1212):

use a calculator to compute \[8^3\] then put it in the denominator

OpenStudy (anonymous):

ok...i got 512 when i multiplied that so now what? @misty1212

OpenStudy (anonymous):

oh so is my answer 1/512 @misty1212

OpenStudy (misty1212):

yes dear that is right

OpenStudy (anonymous):

thank you! ^_^